OpenStudy (anonymous):

write an equation for this relation: {(1,2), (2,4), (3,8), (4,16), (5,32) *this is algebra 1problem

myininaya (myininaya):

\[2^1\] ... \[2^3\] ...

OpenStudy (anonymous):

umm it has to be in a equation

myininaya (myininaya):

i'm showing you a pattern

myininaya (myininaya):

you need to realize the pattern before you can make an equation

myininaya (myininaya):

these are the points you gave me \[(1,2^1);(2,2^2);(3,2^3);(4,2^4);(5,2^5)\] what do they have in common?

OpenStudy (anonymous):

okay so far i see that the patter is increasing the exponent...

myininaya (myininaya):

\[y=2^x\]

OpenStudy (anonymous):

well... i gave you (1,2), (2,4), (3,8), (4,16), (5,32) these...

OpenStudy (anonymous):

how can it be y=2^x?

myininaya (myininaya):

yes thats right you gave me: \[(1,2^1);(2,2^2);(3,2^3);(4,2^4);(5,2^5) \]

myininaya (myininaya):

x=1 => y=2^1=2 x=2 => y=2^2=4 x=3 => y=2^3=8 x=4 => y=2^4=16 x=5 => y=2^5=32

OpenStudy (anonymous):

oops... sorry... i didnt realize it

OpenStudy (anonymous):

so its y=2^x?

OpenStudy (anonymous):

thanks so muchh!!!